OpenStudy (anonymous):

the width of a rectangle is 3 less than twice its width its perimeter is 24 How d i translate this to an equation and find the answer?

OpenStudy (anonymous):

you said width twice in the beginning is one of them supposed to be length

OpenStudy (anonymous):

Yeah Sorry haha

OpenStudy (anonymous):

if it's width is 3 less than twice it's LENGTH, then width = 2L-3, and length is just L. so it'd be 4L- 6+ 2L = 24, so 6L -6 = 24, add 6 to both sides, 6L = 30, Length = 5.

OpenStudy (anonymous):

if that's what you were looking for

OpenStudy (anonymous):

Yeah where did you get the 4L from?

OpenStudy (anonymous):

I multipled 2L-3 by two because there's two sides that are the width, and same for L, because there's two sides that are length.

OpenStudy (anonymous):

Okay. But if the length is 5 that means the width is 7? That doesnt make much sense.

OpenStudy (anonymous):

maybe it's the other way around... length = 7 and width = 5?

OpenStudy (anonymous):

because i was originally confused about how you worded the problem

OpenStudy (anonymous):

Maybe. But the area has t equal 65. I tried L=10 and W=6.5 but it's still nt right

OpenStudy (anonymous):

P = 2L + 2W 24 = 2L + 2(2L - 3) 24 = 2L + 4L - 6 24 = 6L - 6 24 + 6 = 6L 30 = 6L 5 = L W = 2L - 3 W = 2(5) - 3 W = 10 - 3 W = 7 check... P = 2L + 2W 24 = 2(5) + 2(7) 24 = 10 + 14 24 = 24 (correct)

OpenStudy (anonymous):

see she did it the same way but it does NOT make sense that the widh is more than the length

OpenStudy (anonymous):

might have to ask your teacher about this one

OpenStudy (anonymous):

Yes you're right. It's a bit confusing.

OpenStudy (anonymous):

the width is 3 less then twice its length. width is 7. three less then twice its length, means 3 less the 10, which is 7.

OpenStudy (anonymous):

typo * then

OpenStudy (anonymous):

Ohhhhh! I see how now! Thank you both so much! You saved my homework! :)

OpenStudy (anonymous):

you would have eventually got it...you seem like you have pretty good sense.

OpenStudy (anonymous):

Thank you! Have a good night!

OpenStudy (anonymous):

you too :)